OpenStudy (anonymous):

can someone check my answer? -4 3/5 - 2 1/2 -4 6/10 - 2 5/10 -6 9/10

OpenStudy (anonymous):

Actually, it's 7 1/10

OpenStudy (anonymous):

-7 1/10

OpenStudy (anonymous):

you are super close. the answer would be -5 9/10 because ultimately the problem is -34/10 - 25/10 -32-25=-59 and its all over ten. once you put it in proper form its -5 9/10

OpenStudy (anonymous):

@micahwood50 could you show how you got that answer please?

OpenStudy (anonymous):

@abiii818 It's not. Check this: http://www.wolframalpha.com/input/?i=-4%203%2F5%20-%202%201%2F2&t=crmtb01

OpenStudy (anonymous):

ok my bad!

OpenStudy (anonymous):

i did it in my calculator and that was what i got:/

OpenStudy (anonymous):

i put the wrong sign in. it is -7 1/10. my bad!!

OpenStudy (anonymous):

\[-4\frac{3}{5} - 2\frac{1}{2}\] \[-4\frac{3 \cdot 2}{5 \cdot 2} - 2\frac{1 \cdot 5}{2 \cdot 5}\] \[-4\frac{6}{10} - 2\frac{5}{10}\] \[-1\left(4\frac{6}{10} + 2\frac{5}{10}\right)\] \[-1\left(7\frac{1}{10}\right)\] \[-7\frac{1}{10}\]

OpenStudy (anonymous):

thanks so much!!!!!!!!!!!!!!!!!!!!! : D

OpenStudy (anonymous):

Note that \[4\frac{6}{10} + 2\frac{5}{10} = 6\frac{11}{10} = 7\frac{1}{10}\]

OpenStudy (anonymous):

Glad I helped.